OpenStudy (anonymous):

A number is divided by 2. If 16 is added to the quotient you get 20. What is the number? A.2 B.8 C.18 D.72

OpenStudy (anonymous):

If we let n represent the number, what does dividing it by 2 look like?

OpenStudy (anonymous):

20-16=4 \[\frac{ n }{ 2 }=4\]\[2(\frac{ n }{ 2 })=(4)2\]\[n=8\]